OpenStudy (anonymous):

Solve 7x – 2 = 12x Answer x ≈ 3.28 x ≈ 0.31 x ≈ –7.22 x ≈ –0.14

OpenStudy (roadjester):

Isolate the x's on one side and everything without x on the other

OpenStudy (anonymous):

7^x-2=12^x

OpenStudy (roadjester):

Subtract 7x from both sides.\[7x-2=12x\] \[-2=5x\] Divide by 5 \[x={-2\over 5}\] and @josettestuff is correct, -2/5 is -0.4 as a decimal

OpenStudy (roadjester):

That makes things more fun.

OpenStudy (roadjester):

Ok so the question is \[\huge 7^x-2=12^x\]?

OpenStudy (roadjester):

This question cannot be solved as is...It has to be 7^(x-2)=12^x. If that's the case, the answer would be -7.22. You need to put the "^" and parentheses or else people won't understand the question.